# Open-Set Domain Adaptation Under Background Distribution Shift: Challenges and A Provably Efficient Solution

### v2

@@ 4 Constrained Open-Set Learning Rules and Domain Discriminators
We focus on the case $k=1$, i.e., PU-learning (bekker2020learning), to study the novelty detection aspect of OSDA. The domain discrimination approach trains a classifier, via Empirical Risk Minimization (ERM), to distinguish between $P_{{\mathcal{S}}}$ and $P_{{\mathcal{T}}}$ such that for a 0-1 loss function $l_{01}$, $h_{\mathrm{\text{DD}}}=\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\frac{1}{N_{\mathcal{S}}+N_{\mathcal{T}}}\big(\sum_{{\mathbf{x}}\in S_{{\mathcal{S}}}}l_{01}(h({\mathbf{x}}),0)+\sum_{{\mathbf{x}}\in S_{{\mathcal{T}}}}l_{01}(h({\mathbf{x}}),1)\big)$. Similarly, the learning rule for constrained learning is $h_{\mathrm{\text{color}}}=\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\frac{1}{N_{\mathcal{S}}}\big(\sum_{{\mathbf{x}}\in S_{{\mathcal{S}}}}l_{01}(h({\mathbf{x}}),0)\big)\text{ s.t. }\frac{1}{N_{\mathcal{T}}}\sum_{{\mathbf{x}}\in S_{{\mathcal{T}}}}l_{01}(h({\mathbf{x}}),1)\leq 1-\hat{\alpha}$. In practice, these rules are often implemented with the logistic loss or some close variation, e.g. (kiryo2017positive).
As noted in section˜2, standard PU-learning methods without distribution shift rely on domain discriminators. Constrained learning rules were studied in blanchard2010semi and later extended under distribution shift by wald2023birds using small models. Their analysis, however, is limited to classical learning regimes that assume infinite data, whereas in practice, we prefer high-capacity models and only have limited data. A regime that is beyond the scope of their theory. In this section, we review both approaches and explain why applying these guarantees to such high-capacity models is non-trivial. Our main result in section˜4.2 compares constrained learning and domain discrimination in a simplified overparameterized setting that preserves key aspects of open-set training. The model builds on ideas from generalization to minority subgroups (sagawa2020investigation; wald2022malign; puli2023don; nagarajan2021understanding). Empirical results in section˜6 corroborate conclusions to which the theory points.
@@ 4.1 Overparameterization and Finite Sample Analysis
In terms of formal guaranties, it is well known that at the limit of infinite data and a family $\mathcal{H}$} that is large enough, the minimizer $h_{\mathrm{\text{DD}}}({\mathbf{x}})$ will be proportional to the log-odds ratio, $h_{\mathrm{\text{DD}}}({\mathbf{x}})=\log(P_{{\mathcal{T}}}({\mathbf{x}})/P_{{\mathcal{S}}}({\mathbf{x}}))$. Under assumption˜1, this means that the instances of the novel class will be classified as originating from $P_{{\mathcal{T}}}$ w.p. $1$, while for ${\mathbf{x}}\sim P_{{\mathcal{T}},0}$ ($P_{{\mathcal{T}},[k]}$ since $k=1$) it will be bounded away from $1$ since $\mathrm{Supp}(P_{{\mathcal{T}},[k]})\subseteq\mathrm{Supp}(P_{{\mathcal{S}}})$. Hence, under these favorable conditions, $h_{\mathrm{\text{DD}}}$ will classify the novel class vs. the single known class with accuracy $1$. In practice, we do not perform perfect optimization with infinitely large datasets and highly expressive models. Therefore, it is desirable to study the sample complexity of different learning rules. Without further assumptions, the sample complexity of the domain discrimination approach resembles those in the domain adaptation literature, which depend on the divergence between $P_{{\mathcal{S}}}$ and $P_{{\mathcal{T}}}$, e.g. (ben2010theory). wald2023birds show that instead of following ERM, solving a constrained learning rule, achieves a generalization bound where the error scales with a divergence $d_{{\mathcal{H}},\beta}(P_{{\mathcal{S}}}\|P_{{\mathcal{T}}})$ which is usually considerably smaller than common divergences in domain adaptation  [^4] Intuitively, $\beta$ is a small number, and $d_{{\mathcal{H}},\beta}(P_{{\mathcal{S}}}\|P_{{\mathcal{T}}})$ measures how probable a rare event under $P_{{\mathcal{S}}}$, i.e. with probability smaller than $\beta$, can become under $P_{{\mathcal{T}}}$. Here, $\beta$ is a constant that depends on the Rademacher complexity of $\mathcal{H}$} and the sample size. We refer to Thm 4.3 in wald2023birds for more details.
Why existing theory may not reflect large-scale training. As reflected by dependence on quantities like the Rademacher complexity of $\mathcal{H}$}, the results on constrained learning apply to a regime where models are of relatively low capacity and may not be expressive enough to achieve arbitarily low training loss. However, it is rather common to fit much more expressive models, as they often generalize better in practice. In what follows we formally study a simple example of overparameterized models, where the number of parameters is larger than the number of training examples. This is a common toy model to study deep networks that overfit their training data but still generalize well, also known as “benign overfitting” (belkin2019reconciling).
@@
To capture key aspects of the problem solved in practice, while maintaining a manageable mathematical analysis, we perform a few simplifications: Data and models. (i) We study linear models $h({\mathbf{x}})={\mathbf{w}}^{\top}{\mathbf{x}}$ for ${\mathbf{w}}\in{{\mathbb{R}}^{d}}$, where $d>N_{{\mathcal{S}}}+N_{{\mathcal{T}}}$ so the models are expressive and can interpolate the data; (ii) We focus on a problem with two features (along $\mu$ and $\eta$) and Gaussian noise (see definition˜2). Here $\mu$ and $\eta$ denote the two informative directions in the problem: the background shift, i.e. the shift in the known class, is in the direction $\mu$, and its magnitude is proportional to the norm $r_{\mu}$. While $\eta$ marks the novel class, and higher $r_{\eta}$ corresponds to a larger signal for learning the novel class. The covariance matrices (e.g., $I_{d}-r_{\mu}^{-2}\mu\mu^{\top}$) remove the corresponding rank-one components so the problem satisfies assumption˜1. This construction keeps the model analytically tractable and this type of simplifications are common in works on benign overfitting and robustness to distribution shifts (nagarajan2021understanding; wald2022malign; muthukumar2021classification).
@@
Let $d>0$, $\mu,\eta\in{{\mathbb{R}}^{d}}$, $\sigma=1/\sqrt{d}$, $\alpha\in{(0,1)}$, $r_{\mu}=\|\mu\|$, $r_{\eta}=\|\eta\|$. A Linear-Gaussian PU-learning problem is an OSDA problem (definition˜1) with $1$ known class where $P_{{\mathcal{S}}}={\mathcal{N}}(\mu,\sigma^{2}(I_{d}-r_{\eta}^{-2}\eta\eta^{\top}))$, $P_{\mathcal{T},0}={\mathcal{N}}(-\mu,\sigma^{2}(I_{d}-r_{\eta}^{-2}\eta\eta^{\top})$ and $P_{\mathcal{T},1}={\mathcal{N}}(-\eta,\sigma^{2}(I_{d}-r^{-2}_{\mu}\mu\mu^{\top}))$.
Simplified CoLOR and Domain Discriminator methods. (iii) The last simplification regards which optimization problem to analyze. In an overparameterized problem, there are many possible minimizers of the empirical risk and the constrained risk in eq.˜3. Therefore to analyze the solutions, it is common to consider $h_{\mathrm{\text{DD}}}$ and $h_{\mathrm{\text{color}}}$ as max-margin classifiers  [^5] This is based on results that show the implicit bias of gradient descent and the logistic loss to the max-margin classifier on separable data (soudry2018implicit) such that:
@@
Consider a Linear-Gaussian PU-learning problem with parameters $\mu,\eta,d>300,\alpha\in{(0,\frac{1}{1024})}$ and dataset sizes $_{{\mathcal{T}}}>10,N_{{\mathcal{S}}}=N_{{\mathcal{T}}}, N=N$$N_{{\mathcal{S}}}+N_{{\mathcal{T}}}=2N_{{\mathcal{T}}}$, and let $\delta\in{(0,1)}$. For all problems where
$\min{\{r_{\eta},r_{\mu}\}}\geq\frac{32}{\sqrt{N}},r_{\mu}\leq\frac{1}{2}\sqrt{N_{{\mathcal{T}},0}},\frac{r_{\eta}}{r_{\mu}}\leq\frac{32}{N},d>c_{1}N\log\left(\frac{c_{2}}{\delta}\right),N\geq c_{1}\log(\frac{c_{2}}{\delta}),$
@@
The proof consists of four parts. First, we show that $\mathrm{AU-ROC}({\mathbf{w}})=Q(\frac{{\mathbf{w}}^{\top}(\mu-\eta)}{\sigma\sqrt{\|{\mathbf{w}}\|^{2}-({\mathbf{w}}^{\top}(\mu+\eta))^{2}}})$, where $Q(\cdot)$ is the inverse Gaussian tail function. This means that whenever ${\mathbf{w}}_{\mathrm{\text{DD}}}^{\top}(\mu-\eta)\leq 0$, the domain discriminator algorithm will have $\mathrm{AU-ROC}({\mathbf{w}}_{\mathrm{\text{DD}}})\leq 0.5$. To derive ranges of problem parameters where this happens, we analyze the convex dual of the maximum-margin problem with an added constraint that ${\mathbf{w}}_{\mathrm{\text{DD}}}^{\top}(\mu-\eta)>0$ and obtain a lower bound $\underline{\gamma}$ on the value of the primal problem (i.e. a lower bound on the norm of a “good” solution). Then we guess a solution of the form ${\mathbf{w}}^{\prime}=\mu+\sum_{i\in{U}}{\xi_{i}}$, which does not depend on $\eta$ at all and has AU-ROC smaller than $0.5$. Intuitively, $U$ is constructed by including the novel class examples and use the noise vectors of these examples to fit them and satisfy the constraints of the problem. Then, by showing that $\|{\mathbf{w}}^{\prime}\|<\underline{\gamma}$ we prove that $\mathrm{AU-ROC}({\mathbf{w}}_{\mathrm{\text{DD}}})\leq 0.5$. This is summarized in proposition˜1. Next in proposition˜2, we follow a similar procedure with ${\mathbf{w}}_{\mathrm{\text{color}}}$ of analyzing a problem with an added constraint ${\mathbf{w}}^{\top}(\mu-\eta)<\tau$ for some $\tau>0$. Again using weak duality and guessing a solution, we prove that ${\mathbf{w}}_{\mathrm{\text{color}}}$ will satisfy the constraint, this time showing that it obtains a high AU-ROC. Finally, by finding an intersection of parameters ranges where both bounds on the AU-ROC hold, we prove the main claim in theorem˜2 (restatement of theorem˜1). ∎
@@ 5 CoLOR: A solution to OSDA under Background Shift
In practice, we formulate constrained learning rule (CoLOR) for OSDA setup under background shift as follows:
$h_{\text{color}}=\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\!\left(\sum_{i\in S_{{\mathcal{S}}}}l_{ce}(h_{c}({\mathbf{x}}_{i}),y_{i})+\hat{\beta}(h_{\mathrm{\text{novel}}})\right)\ \text{s.t.}\ \hat{\alpha}(h_{\mathrm{\text{novel}}})<\hat{\alpha}$. (3)
where $h=[h_{c},h_{\mathrm{\text{novel}}}]$, $h_{c}$ are classification heads for known classes, $h_{\mathrm{\text{novel}}}$ is the novelty detection head, $l_{ce}$ is the supervised cross entropy loss, $\hat{\alpha}(h),\hat{\beta}(h)$ are empirical estimates of $\alpha(h),\beta(h)$ from $S_{{\mathcal{T}}}$ and $S_{{\mathcal{S}}}$ respectively. $\hat{\alpha}$ is some constant corresponding to the target recall. It is a hyperparameter that needs tuning.
@@ 5.1 Efficient Architecture for Estimating the Novel Class Ratio ($\alpha$)
To solve the eq.˜3, we follow wald2023birds; chamon2022constrained; cotter2019optimization and solve a Lagrangian optimization problem obtained by switching the role of maximization and constraints in eq.˜3 and taking its dual, see eq.˜4 for the full objective. However, a naïve implementation of this objective has a significant drawback in terms of computational complexity due to hyperparameter optimization. Indeed, prior work on the rate-constrained learning problems we seek to solve (where rate in our problem corresponds to the size of the novel class) is limited to either very small models and datasets (wald2023birds), or to certain applications such as fairness (zafar2019fairness), where the desired rate is known a-priori.
The computationally challenging part is that our Lagrangian problem needs to be solved for $L$ candidate values of $\hat{\alpha}\in{\bm{\alpha}}$, corresponding to the potential estimates of novel class proportion, where $\bm{\alpha}\in{[0,1]^{L}}$. We perform grid search over $\bm{\alpha}$ and choose the model that obtains largest empirical recall estimate ($\hat{\alpha}(h)$) w.r.t. the task of distinguishing between $P_{{\mathcal{S}}}$ and $P_{{\mathcal{T}}}$, while still satisfying the constraint on empirical FPR i.e. false positive rate estimate ($\hat{\beta}$) for the same task as specified in the learning objective. For empirical purposes, we calculate the approximate bound for $\beta$ (constraint on $\hat{\beta}(h)$) using the Rademacher complexity in theorem 1 of wald2023birds. We find that setting $\beta=0.01$ is well within the theoretically calculated bounds and works well in practice for all the experiments. Plots in figure 4 provide further insights about the impact of varying $\beta$ on OSDA performance of CoLOR.
@@
Figure 2: (left) CoLOR architecture for OSDA, heads $w^{a}_{i}$ for multiple recall values and classification heads $w^{c}$, vs. (right) a network optimizing for novelty detection with single recall value as in wald2023birds.

### v3

@@ 4 Constrained Open-Set Learning Rules and Domain Discriminators
We focus on the case $k=1$, i.e., PU-learning (Bekker and Davis, 2020), to study the novelty detection aspect of OSDA. The domain discrimination approach trains a classifier, via Empirical Risk Minimization (ERM), to distinguish between $P_{{\mathcal{S}}}$ and $P_{{\mathcal{T}}}$ such that for a 0-1 loss function $l_{01}$, $h_{\mathrm{\text{DD}}}=\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\frac{1}{N_{\mathcal{S}}+N_{\mathcal{T}}}\big(\sum_{{\mathbf{x}}\in S_{{\mathcal{S}}}}l_{01}(h({\mathbf{x}}),0)+\sum_{{\mathbf{x}}\in S_{{\mathcal{T}}}}l_{01}(h({\mathbf{x}}),1)\big)$. Similarly, the learning rule for constrained learning is $h_{\mathrm{\text{color}}}=\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\frac{1}{N_{\mathcal{S}}}\big(\sum_{{\mathbf{x}}\in S_{{\mathcal{S}}}}l_{01}(h({\mathbf{x}}),0)\big)\text{ s.t. }\frac{1}{N_{\mathcal{T}}}\sum_{{\mathbf{x}}\in S_{{\mathcal{T}}}}l_{01}(h({\mathbf{x}}),1)\leq 1-\hat{\alpha}$. In practice, these rules are often implemented with the logistic loss or some close variation, e.g. (Kiryo et al., 2017).
As noted in section˜2, standard PU-learning methods without distribution shift rely on domain discriminators. Constrained learning rules were studied in Blanchard et al. (2010) and later extended under distribution shift by Wald and Saria (2023) using small models. Their analysis, however, is limited to classical learning regimes that assume infinite data, whereas in practice, we prefer high-capacity models and only have limited data. A regime that is beyond the scope of their theory. In this section, we review both approaches and explain why applying these guarantees to such high-capacity models is non-trivial. Our main result in section˜4.2 compares constrained learning and domain discrimination in a simplified overparameterized setting that preserves key aspects of open-set training. The model builds on ideas from generalization to minority subgroups (Sagawa et al., 2020; Wald et al., 2022; Puli et al., 2023; Nagarajan et al., 2021). Empirical results in section˜6 corroborate conclusions to which the theory points.
@@ 4.1 Overparameterization and Finite Sample Analysis
In terms of formal guaranties, it is well known that at the limit of infinite data and a family $\mathcal{H}$} that is large enough, the minimizer $h_{\mathrm{\text{DD}}}({\mathbf{x}})$ will be proportional to the log-odds ratio, $h_{\mathrm{\text{DD}}}({\mathbf{x}})=\log(P_{{\mathcal{T}}}({\mathbf{x}})/P_{{\mathcal{S}}}({\mathbf{x}}))$. Under assumption˜1, this means that the instances of the novel class will be classified as originating from $P_{{\mathcal{T}}}$ w.p. $1$, while for ${\mathbf{x}}\sim P_{{\mathcal{T}},0}$ ($P_{{\mathcal{T}},[k]}$ since $k=1$) it will be bounded away from $1$ since $\mathrm{Supp}(P_{{\mathcal{T}},[k]})\subseteq\mathrm{Supp}(P_{{\mathcal{S}}})$. Hence, under these favorable conditions, $h_{\mathrm{\text{DD}}}$ will classify the novel class vs. the single known class with accuracy $1$. In practice, we do not perform perfect optimization with infinitely large datasets and highly expressive models. Therefore, it is desirable to study the sample complexity of different learning rules. Without further assumptions, the sample complexity of the domain discrimination approach resembles those in the domain adaptation literature, which depend on the divergence between $P_{{\mathcal{S}}}$ and $P_{{\mathcal{T}}}$, e.g. (Ben-David et al., 2010a). Wald and Saria (2023) show that instead of following ERM, solving a constrained learning rule, achieves a generalization bound where the error scales with a divergence $d_{{\mathcal{H}},\beta}(P_{{\mathcal{S}}}\|P_{{\mathcal{T}}})$ which is usually considerably smaller than common divergences in domain adaptation  [^4] Intuitively, $\beta$ is a small number, and $d_{{\mathcal{H}},\beta}(P_{{\mathcal{S}}}\|P_{{\mathcal{T}}})$ measures how probable a rare event under $P_{{\mathcal{S}}}$, i.e. with probability smaller than $\beta$, can become under $P_{{\mathcal{T}}}$. Here, $\beta$ is a constant that depends on the Rademacher complexity of $\mathcal{H}$} and the sample size. We refer to Thm 4.3 in Wald and Saria (2023) for more details.
Why existing theory may not reflect large-scale training. As reflected by dependence on quantities like the Rademacher complexity of $\mathcal{H}$}, the results on constrained learning apply to a regime where models are of relatively low capacity and may not be expressive enough to achieve arbitarily low training loss. However, it is rather common to fit much more expressive models, as they often generalize better in practice. In what follows we formally study a simple example of overparameterized models, where the number of parameters is larger than the number of training examples. This is a common toy model to study deep networks that overfit their training data but still generalize well, also known as “benign overfitting” (Belkin et al., 2019).
@@
To capture key aspects of the problem solved in practice, while maintaining a manageable mathematical analysis, we perform a few simplifications: Data and models. (i) We study linear models $h({\mathbf{x}})={\mathbf{w}}^{\top}{\mathbf{x}}$ for ${\mathbf{w}}\in{{\mathbb{R}}^{d}}$, where $d>N_{{\mathcal{S}}}+N_{{\mathcal{T}}}$ so the models are expressive and can interpolate the data; (ii) We focus on a problem with two features (along $\bm{\mu}$ and $\bm{\eta}$) and Gaussian noise (see definition˜2). Here $\bm{\mu}$ and $\bm{\eta}$ denote the two informative directions in the problem: the background shift, i.e. the shift in the known class, is in the direction $\bm{\mu}$, and its magnitude is proportional to the norm $r_{\mu}$. While $\bm{\eta}$ marks the novel class, and higher $r_{\eta}$ corresponds to a larger signal for learning the novel class. The covariance matrices (e.g., $I_{d}-r_{\mu}^{-2}\bm{\mu}\bm{\mu}^{\top}$) remove the corresponding rank-one components so the problem satisfies assumption˜1. This construction keeps the model analytically tractable and this type of simplifications are common in works on benign overfitting and robustness to distribution shifts (Nagarajan et al., 2021; Wald et al., 2022; Muthukumar et al., 2021).
@@
Let $d>0$, $\bm{\mu},\bm{\eta}\in{{\mathbb{R}}^{d}}$, $\sigma=1/\sqrt{d}$, $\alpha\in{(0,1)}$, $r_{\mu}=\|\bm{\mu}\|$, $r_{\eta}=\|\bm{\eta}\|$. A Linear-Gaussian PU-learning problem is an OSDA problem (definition˜1) with $1$ known class where $P_{{\mathcal{S}}}={\mathcal{N}}(\bm{\mu},\sigma^{2}(I_{d}-r_{\eta}^{-2}\bm{\eta}\bm{\eta}^{\top}))$, $P_{\mathcal{T},0}={\mathcal{N}}(-\bm{\mu},\sigma^{2}(I_{d}-r_{\eta}^{-2}\bm{\eta}\bm{\eta}^{\top})$ and $P_{\mathcal{T},1}={\mathcal{N}}(-\bm{\eta},\sigma^{2}(I_{d}-r^{-2}_{\mu}\bm{\mu}\bm{\mu}^{\top}))$.
Simplified CoLOR and Domain Discriminator methods. (iii) The last simplification regards which optimization problem to analyze. In an overparameterized problem, there are many possible minimizers of the empirical risk and the constrained risk in eq.˜3. Therefore to analyze the solutions, it is common to consider $h_{\mathrm{\text{DD}}}$ and $h_{\mathrm{\text{color}}}$ as max-margin classifiers  [^5] This is based on results that show the implicit bias of gradient descent and the logistic loss to the max-margin classifier on separable data (Soudry et al., 2018) such that:
@@
Consider a Linear-Gaussian PU-learning problem with parameters $\bm{\mu},\bm{\eta},d>300$, dataset sizes $_{{\mathcal{T}}}>10, N=N$$N_{{\mathcal{S}}}+N_{{\mathcal{T}}}$, $\alpha\in{(0,\frac{N}{1024N_{\mathcal{T}}})}$ and let $\delta\in{(0,1)}$. For all problems where
$\min\{r_{\eta},r_{\mu}\}\geq\tfrac{16}{\sqrt{N_{\mathcal{T}}}},\quad r_{\mu}\leq\tfrac{1}{2}\sqrt{N_{{\mathcal{T}},0}},\quad\tfrac{r_{\eta}}{r_{\mu}}\leq\tfrac{4}{N},\quad c_{1}\log\!\left(\tfrac{c_{2}}{\delta}\right)\leq\min{\left(\sqrt{N},\sqrt{\frac{d}{N}}\right)},$
@@
The proof consists of four parts. First, we show that $\mathrm{AU-ROC}({\mathbf{w}})=Q(\frac{{\mathbf{w}}^{\top}(\bm{\mu}-\bm{\eta})}{\sigma\sqrt{\|{\mathbf{w}}\|^{2}-({\mathbf{w}}^{\top}(\bm{\mu}+\bm{\eta}))^{2}}})$, where $Q(\cdot)$ is the inverse Gaussian tail function. This means that whenever ${\mathbf{w}}_{\mathrm{\text{DD}}}^{\top}(\bm{\mu}-\bm{\eta})\leq 0$, the domain discriminator algorithm will have $\mathrm{AU-ROC}({\mathbf{w}}_{\mathrm{\text{DD}}})\leq 0.5$. To derive ranges of problem parameters where this happens, we analyze the convex dual of the maximum-margin problem with an added constraint that ${\mathbf{w}}_{\mathrm{\text{DD}}}^{\top}(\bm{\mu}-\bm{\eta})>0$ and obtain a lower bound $\underline{\gamma}$ on the value of the primal problem (i.e. a lower bound on the norm of a “good” solution). Then we guess a solution of the form ${\mathbf{w}}^{\prime}=\bm{\mu}+\sum_{i\in{U}}{\xi_{i}}$, which does not depend on $\bm{\eta}$ at all and has AU-ROC smaller than $0.5$. Intuitively, $U$ is constructed by including the novel class examples and use the noise vectors of these examples to fit them and satisfy the constraints of the problem. Then, by showing that $\|{\mathbf{w}}^{\prime}\|<\underline{\gamma}$ we prove that $\mathrm{AU-ROC}({\mathbf{w}}_{\mathrm{\text{DD}}})\leq 0.5$. This is summarized in proposition˜1. Next in proposition˜2, we follow a similar procedure with ${\mathbf{w}}_{\mathrm{\text{color}}}$ of analyzing a problem with an added constraint ${\mathbf{w}}^{\top}(\bm{\mu}-\bm{\eta})<\tau$ for some $\tau>0$. Again using weak duality and guessing a solution, we prove that ${\mathbf{w}}_{\mathrm{\text{color}}}$ will satisfy the constraint, this time showing that it obtains a high AU-ROC. Finally, by finding an intersection of parameters ranges where both bounds on the AU-ROC hold, we prove the main claim in theorem˜2 (restatement of theorem˜1). ∎
@@ 5 CoLOR: A solution to OSDA under Background Shift
Below we formulate constrained learning rule (CoLOR) for OSDA setup under background shift for a true novel class proportion ($\alpha$) in the target dataset and some tolerance $\beta>0$ for misclassification of $S_{\mathcal{S}}$ by $h_{\mathrm{\text{novel}}}$:
$h_{\text{color}}=\operatorname*{arg\,min}_{h\in{\mathcal{H}}}\!\left(\sum_{i\in S_{{\mathcal{S}}}}l_{ce}(h_{c}({\mathbf{x}}_{i}),y_{i})+\hat{\beta}_{\mathrm{emp}}(h_{\mathrm{\text{novel}}})\right)\ \text{s.t.}\ \hat{\alpha}_{\mathrm{emp}}(h_{\mathrm{\text{novel}}})<\hat{\alpha}$. (3)
where $h=[h_{c},h_{\mathrm{\text{novel}}}]$, $h_{c}$ are classification heads for known classes, $h_{\mathrm{\text{novel}}}$ is the novelty detection head, $l_{ce}$ is the supervised cross entropy loss, $\hat{\alpha}_{\mathrm{emp}}(h),\hat{\beta}_{\mathrm{emp}}(h)$ are empirical estimates of $\alpha(h),\beta(h)$ over $S_{{\mathcal{T}}}$ and $S_{{\mathcal{S}}}$ respectively. $\hat{\alpha}$ is some constant corresponding to the target recall. It is a hyperparameter that needs tuning.
@@ 5.1 Efficient Architecture for Estimating the Novel Class Ratio ($\alpha$)
To solve the eq.˜3, we follow Wald and Saria (2023); Chamon et al. (2022); Cotter et al. (2019) and solve a Lagrangian optimization problem obtained by switching the role of maximization and constraints in eq.˜3 and taking its dual, see eq.˜4 for the full objective. However, a naïve implementation of this objective has a significant drawback in terms of computational complexity due to hyperparameter optimization. Indeed, prior work on the rate-constrained learning problems we seek to solve (where rate in our problem corresponds to the size of the novel class) is limited to either very small models and datasets (Wald and Saria, 2023), or to certain applications such as fairness (Zafar et al., 2019), where the desired rate is known a-priori.
The computationally challenging part is that our Lagrangian problem needs to be solved for $L$ candidate values of $\hat{\alpha}\in{\bm{\alpha}}$, corresponding to the potential estimates of novel class proportion, where $\bm{\alpha}\in{[0,1]^{L}}$. We perform grid search over $\bm{\alpha}$ and choose the model that obtains largest empirical recall estimate ($\hat{\alpha}(h)$) w.r.t. the task of distinguishing between $P_{{\mathcal{S}}}$ and $P_{{\mathcal{T}}}$, while still satisfying the constraint on empirical FPR i.e. false positive rate estimate ($\hat{\beta}$) for the same task as specified in the learning objective. For empirical purposes, we calculate the approximate bound for $\beta$ (constraint on $\hat{\beta}(h)$) using the Rademacher complexity in theorem 1 of Wald and Saria (2023). We find that setting $\beta=0.01$ is well within the theoretically calculated bounds and works well in practice for all the experiments. Plots in figure 4 provide further insights about the impact of varying $\beta$ on OSDA performance of CoLOR.
@@
Figure 2: (left) CoLOR architecture for OSDA, heads $w^{a}_{i}$ for multiple recall values and classification heads $w^{c}$, vs. (right) a network optimizing for novelty detection with single recall value as in Wald and Saria (2023).